# A differentiated replication study on LLMs

### v1

@@ V-3 Models
  • M1: Models should have already been evaluated in similar pipeline. Both the evaluations of LLMs for AVR performed by Wu et al. [5] and Pearce et al. [9] respectively found that Codex [33] was the best performing model. Unfortunately, Codex models are no longer available, so Kulsum et al. [7] had to base their approach on ChatGPT 3.5 [34], We build on their choice and adopt it as the first model in our study.
  • M2: Best proprietary model and best open source model on LiveBench [35] At the time of writing, the best-performing proprietary model is GPT-o4-Mini High [36] and the best open-source model is DeepSeek R1 Distill Qwen 32B [37].
@@ VI-2 Error introduction
We introduce a controlled error in vulnerability localization. As in Setup, each LLM receives the full vulnerable function as input. Since correction is impossible when localization is incorrect at the function level, we apply this step only when line-level localization is also provided. For each such prompt, we generate six variants by shifting the marked line by 2, 4, 8 lines—either above (e.g., -2) or below (e.g., +2) the correct location. Variants that fall outside the function’s scope are discarded. If testing all prompts becomes too costly, we apply a Taguchi design [38, 39] to design a balanced experiment and reduce the number of evaluations.
@@ VI-3 Patch generation
We query each LLM (or LLM-based setup, when we replicate VRPilot [7]) with all prompts from §V-2, or their error-injected versions. Each model is queried exactly once per prompt version and tasked with generating a repaired version of the vulnerable function. The proposed function is then extracted and passed to the validation process.
@@ VI-6 Manual inspection
To answer RQ3, we perform a manual inspection of a random sample of the patches that pass the testing phase. Two authors will independently go through each patch in the sample and assess the correctness of the patch. We refer to Bui et al. [3] and use “Correct” (fix the vulnerability and do not break the functionality), “Partially correct” (fix the vulnerability but break the functionality), and “Wrong” (do not fix the vulnerability) for the classification. The two reviewers will then resolve any disagreements through discussion, or possibly with a third individual as adjudicator. We also mark whether the LLM reasoning is right or wrong.
@@ VII-1 RQ1
To verify the hypotheses about the prompt impact that we presented in Section VI, we will measure statistically significant effects on the number of produced patches passing each phase (generation, functional validation, security validation, manual inspection). We use TOST (two one-sided tests) for testing equivalence [40, 27, 28] with Mann-Whitney-U (MWU) as the underlying directional test with Helmert contrast for multiple comparisons [41] in the configuration shown in Figure 3.
@@ VII-2 $RQ1_{ERR}$
As represented in Figure 4, first, we compare the baseline with all the patches generated with errors in the line-level vulnerability localization. Then, the patches with the smallest error (222) against the ones with higher ones (444, 888). We need a TOST to show a significant equivalence, while a single MWU-will be sufficient to prove a significant difference.
@@ VII-4 $RQ3$
The Agresti-Coull-Wilson method [42] allows to establish accurate and reliable confidence intervals for proportions. This interval will enable us to extrapolate the false positive rate from sampled subsets of patches. By using Cochran’s formula we need to analyze 96 patches to have 95% confidence interval with a 10% margin of error [41]
@@ IX Acknowledgements
This work has been partly supported by the European Union (EU) under Horizon Europe grant n. 101120393 (Sec4AI4Sec), by the Italian Ministry of University and Research (MUR), under the P.N.R.R. – NextGenerationEU grant n. PE00000014 (SERICS), and by the Nederlandse Organisatie voor Wetenschappelijk Onderzoek (NWO) under grant n. KIC1.VE01.20.004 (HEWSTI).

### v2

@@ V-3 Models
M1: Models should have already been evaluated in similar pipeline. Both the evaluations of LLMs for AVR performed by Wu et al. [wu2023effective] and Pearce et al. [pearce2023examining] respectively found that Codex [openai2021codex] was the best performing model. Unfortunately, Codex models are no longer available, so Kulsum et al. [kulsum2024case] had to base their approach on ChatGPT 3.5 [openai2022chatgpt], We build on their choice and adopt it as the first model in our study.
M2: Best proprietary model and best open source model on LiveBench [whitelivebench] At the time of writing, the best-performing proprietary model is GPT-o4-Mini High [openai2025o4mini] and the best open-source model is DeepSeek R1 Distill Qwen 32B [guo2025deepseekr1].
@@ VI-2 Error introduction
We introduce a controlled error in vulnerability localization. As in Setup, each LLM receives the full vulnerable function as input. Since correction is impossible when localization is incorrect at the function level, we apply this step only when line-level localization is also provided. For each such prompt, we generate six variants by shifting the marked line by 2, 4, 8 lines—either above (e.g., -2) or below (e.g., +2) the correct location. Variants that fall outside the function’s scope are discarded. If testing all prompts becomes too costly, we apply a Taguchi design [mitra2011taguchi, massacci2024addressing] to design a balanced experiment and reduce the number of evaluations.
@@ VI-3 Patch generation
We query each LLM (or LLM-based setup, when we replicate VRPilot [kulsum2024case]) with all prompts from §V-2, or their error-injected versions. Each model is queried exactly once per prompt version and tasked with generating a repaired version of the vulnerable function. The proposed function is then extracted and passed to the validation process.
@@ VI-6 Manual inspection
To answer RQ3, we perform a manual inspection of a random sample of the patches that pass the testing phase. Two authors will independently go through each patch in the sample and assess the correctness of the patch. We refer to Bui et al. [bui2024apr4vul] and use “Correct” (fix the vulnerability and do not break the functionality), “Partially correct” (fix the vulnerability but break the functionality), and “Wrong” (do not fix the vulnerability) for the classification. The two reviewers will then resolve any disagreements through discussion, or possibly with a third individual as adjudicator. We also mark whether the LLM reasoning is right or wrong.
@@ VII-1 RQ1
To verify the hypotheses about the prompt impact that we presented in Section VI, we will measure statistically significant effects on the number of produced patches passing each phase (generation, functional validation, security validation, manual inspection). We use TOST (two one-sided tests) for testing equivalence [schuirmann1981hypothesis, food2001guidance, meyners2012equivalence] with Mann-Whitney-U (MWU) as the underlying directional test with Helmert contrast for multiple comparisons [spssstatistics] in the configuration shown in Figure 3.
@@ VII-2 $RQ1_{ERR}$
As represented in Figure 4, first, we compare the baseline with all the patches generated with errors in the line-level vulnerability localization. Then, the patches with the smallest error ($2$) against the ones with higher ones ($4$, 88). We need a TOST to show a significant equivalence, while a single MWU-will be sufficient to prove a significant difference.
@@ VII-4 $RQ3$
The Agresti-Coull-Wilson method [agresti1998approximate] allows to establish accurate and reliable confidence intervals for proportions. This interval will enable us to extrapolate the false positive rate from sampled subsets of patches. By using Cochran’s formula we need to analyze 96 patches to have 95% confidence interval with a 10% margin of error [spssstatistics]
@@ IX Acknowledgements
This work has been partly supported by the European Union (EU) under Horizon Europe grant n. 101120393 (Sec4AI4Sec), by the Italian Ministry of University and Research (MUR), under the P.N.R.R. – NextGenerationEU grant n. PE00000014 (SERICS) CUP E63C24000590001, and by the Nederlandse Organisatie voor Wetenschappelijk Onderzoek (NWO) under grant n. KIC1.VE01.20.004 (HEWSTI).